# Self-testing in the compiled setting via tilted-CHSH inequalities

### v1

@@ 4.2 Compiled self-test for tilted-CHSH inequalities
We now show that the extended tilted-CHSH Bell expressions are a compiled self-test according to Eq. 3.21. To do so, we use the fact that the quantum bound is preserved under the compilation procedure (cf. Theorem 4.2), along with the SOS decomposition given by [BSB24] to derive a bound of the form in Eq. 3.21 via a series of triangle inequalities. Our approach follows closely that of Bamps and Pironio [BP15], who obtained robust self-testing bounds for the Bell expression $T_{\theta}$ in the nonlocal scenario, and includes similar calculations to those used in [NZ23, BVB+24], which establish rigidity statements in the compiled setting.
@@
Let $\theta\in(0,\pi/4]$, $\phi\in\big{(}\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big{)}\setminus\{0\}$, and let $I_{\theta,\phi}$ be the generalized tilted-CHSH functional with quantum bound $\eta^{\mathrm{Q}}_{\theta,\phi}$, given by Eq. 4.2. Then the inequality $I_{\theta,\phi}\leq\eta^{\mathrm{Q}}_{\theta,\phi}$ is a compiled self-test according to Eq. 3.21.
The remainder of this section is devoted to proving the above theorem. Recall the SOS decomposition for $\bar{S}_{\theta,\phi}$, given by the polynomials $N_{0},N_{1}$ defined in Eq. 4.3. Then for compiled models $I_{\theta,\phi}=\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\geq\eta^{\mathrm{Q}}_{\theta,\phi}-\epsilon$, for some $\epsilon\geq 0$, implies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]=\eta^{\mathrm{Q}}_{\theta,\phi}-\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\leq\epsilon$. Using the calculation in Eq. 3.15 with $P=N_{0}$ and $P=N_{1}$ respectively, we find
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]$ $\approx_{\delta_{\mathsf{QHE}}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\Big{|}(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\Big{|}^{2}|\Psi_{\alpha|\chi}\rangle=:\hat{E}_{0},$ (4.11)
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}]$ $\approx_{\delta_{\mathsf{QHE}}}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\Big{|}\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)X_{B}-\cos(2\theta)Z_{B}\Big{|}^{2}|\Psi_{\alpha|\chi}\rangle=:\hat{E}_{1},$
where we defined $Z_{B}:=(B_{0}+B_{1})/(2\cos(\phi))$ and $X_{B}:=(B_{0}-B_{1})/(2\sin(\phi))$, and note that $\hat{E}_{i}\geq 0$ for $i=0,1$. This implies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{i}^{\dagger}N_{i}]\geq\hat{E}_{i}-\delta_{\mathsf{QHE}}$, from which we observe that
$\hat{E}_{0}+\lambda^{2}_{\theta,\phi}\hat{E}_{1}\leq\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]+\lambda^{2}_{\theta,\phi}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}]+\delta^{\prime}_{\mathsf{QHE}}=\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]+\delta^{\prime}_{\mathsf{QHE}}\leq\epsilon+\delta^{\prime}_{\mathsf{QHE}},$ (4.12)
where $\delta^{\prime}_{\mathsf{QHE}}=(1+\lambda^{2}_{\theta,\phi})\delta_{\mathsf{QHE}}$. Since $\hat{E}_{i}\geq 0$, we arrive at
$\hat{E}_{0}\leq\epsilon+\delta^{\prime}_{\mathsf{QHE}}\ \ \mathrm{and}\ \ \lambda^{2}_{\theta,\phi}\hat{E}_{1}\leq\epsilon+\delta^{\prime}_{\mathsf{QHE}}.$ (4.13)
@@
Extracting the states: We begin by studying the relations imposed by $\hat{E}_{0}\leq\epsilon+\delta^{\prime}_{\mathsf{QHE}}$. Define $\delta_{0}:=\epsilon+\delta^{\prime}_{\mathsf{QHE}}$. We begin with several claims:
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}\big{(}\mathbbm{1}-(Z_{B})^{2}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}\big{(}1+1/(2\cos^{2}(\phi))\big{)}=:\delta_{1}$.
@@
$\big{\|}\big{(}\mathbbm{1}-(Z_{B})^{2}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $\leq\big{\|}Z_{B}\big{(}(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\ \ \ \ +\big{\|}(-1)^{\mathsf{Dec}(\alpha)}\big{(}(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$\leq(1+\|Z_{B}\|_{\mathrm{op}}^{2})\big{\|}\big{(}(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
In the first inequality, we used the triangle inequality and the bound on $\hat{E}_{0}$, and for the second inequality, we used the definition of the operator norm $\|A|\psi\rangle\|^{2}\leq\|A\|_{\mathrm{op}}^{2}\||\psi\rangle\|^{2},\ \ \forall\;|\psi\rangle\in\mathcal{H}$. By the definition of $Z_{B}$, $\|Z_{B}\|_{\mathrm{op}}^{2}\leq(1/(4\cos^{2}(\phi)))(\|B_{0}\|_{\mathrm{op}}^{2}+\|B_{1}\|_{\mathrm{op}}^{2})=1/(2\cos^{2}(\phi))$. Taking the expectation over $\chi:\mathsf{Enc}(x=0)=\chi$ and summing over $\alpha$, we apply 1 to get the desired bound. ∎
@@
where we used the same steps in 5 for the first equality, and the operator inequality $\mathbbm{1}+|X_{B}|\geq\mathbbm{1}$ for the inequality. Notice that $|X_{B}|^{2}=(X_{B})^{2}$, which satisfies $\cos^{2}(\phi)(Z_{B})^{2}+\sin^{2}(\phi)(X_{B})^{2}=\mathbbm{1}$ by definition. Plugging this in we find that
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}\big{(}P_{B}^{b}-\delta_{\mathsf{Dec}(\alpha),b}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}/2$.
@@
$\|\big{(}P_{B}^{b}-\delta_{\mathsf{Dec}(\alpha),b}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $=\frac{1}{4}\|\big{(}\tilde{Z}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
@@
$\|\big{(}\tilde{Z}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\|\big{(}Z_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}+\|\big{(}\tilde{Z}_{B}-Z_{B}\big{)}|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
and taking the expectation over $\chi:\mathsf{Enc}(x=0)=\chi$, summing over $\alpha$ and applying 1 and 5 we arrive at the bound. ∎
@@
This isometry is based on the “SWAP gate isometry” [MYS12, YN13]. The isometry is efficient according to Definition 2.2 and has several applications in the bipartite setting. Using 7, we establish the following lemma:
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big{\|}V|\Psi_{\alpha|\chi}\rangle-|\mathsf{Dec}(\alpha)\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2}\leq\delta_{0}.$ (4.15)

### v3

@@ 3.3 Quantum bounds for compiled tilted-CHSH expressions
4 Self-testing in the compiled setting
Recall that a bipartite (quantum) model $\mathrm{Q}$, consists of a shared state $|\Psi\rangle$, along with local POVM measurements $\{M_{a|x}\}$ and $\{N_{b|y}\}$ for Alice and Bob, respectively. Given a Bell expression $I$, the inequality $I\leq\eta^{Q}$ self-tests an ideal bipartite model $\mathrm{Q}^{*}$ if any optimal bipartite model is essentially the same as $\mathrm{Q}^{*}$, modulo some physically irrelevant degrees of freedom. This is more formally stated in terms or the existence of local isometries which maps the employed model to the ideal one. When small errors are permitted, one considers the following definition of robust self-testing.
Definition 4.1 (Bipartite self-test).
The inequality $I\leq\eta^{\mathrm{Q}}$ is a self-test for a bipartite model $\mathrm{Q}^{*}=\left(\{P_{a|x}\},\{Q_{b|y}\},|\phi\rangle\right)$ if there exist a non-negative function $f(\epsilon)$ such that $f(\epsilon)\to 0$ as $\epsilon\to 0$, such that for any bipartite model $\mathrm{Q}=\left(\{M_{a|x}\},\{N_{b|y}\},|\Psi\rangle\right)$ achieving $I\geq\eta^{\mathrm{Q}}-\epsilon$ for $\epsilon\geq 0$, there exists a Hilbert space $\mathcal{H}_{\mathrm{aux}}$, an auxiliary state $|\zeta\rangle\in\mathcal{H}_{\mathrm{aux}}$ and local isometries $V_{A}$ and $V_{B}$, such that defining $\mathcal{H}_{A}\otimes\mathcal{H}_{B}\to\mathbb{C}^{d}\otimes\mathbb{C}^{d}\otimes\mathcal{H}_{\mathrm{aux}}, V=V$$V_{A}\otimes V_{B}$, the following is satisfied for all $x,y,a,b$:
$\big\|V_{A}\otimes V_{B}(M_{a|x}\otimes N_{b|y})|\Psi\rangle-(P_{a|x}\otimes Q_{b|y})|\phi\rangle\otimes|\zeta\rangle\big\|\leq f(\epsilon).$
In the bipartite setting, one could consider the situation where Alice measures first using a POVM $\{P_{a|x}\}$, collapsing the state to a post-measurement state $\rho_{a|x}$ on Bob’s subsystem $\mathcal{H}_{B}$, upon which Bob performs his measurement, resulting in the application of the POVM element $Q_{b|y}$. With this in mind, we consider the setting where the only relevant features of the model are those from Bob’s (resp. Alice’s) perspective. In particular, subsystem $A$ is traced out following the recorded measurement of outcome of $a$ given $x$.
Definition 4.2 (Partial model).
Given a bipartite model $\mathrm{Q}=\left(\{M_{a|x}\},\{N_{b|y}\},|\Psi\rangle\right)$, we define the partial model of $\mathrm{Q}$ by $\mathrm{Q}^{\prime}=(\{N_{b|y}\},\{\rho_{a|x}\})$ where
$\rho_{a|x}={\mathrm{t}r}_{A}[(M_{a|x}\otimes\mathbbm{1}_{B})|\Psi\rangle\!\langle\Psi|].$ (4.1)
We note that $\rho_{a|x}$ will generally be mixed. When each $\rho_{a|x}$ is pure, we say that $\mathrm{Q}$ has a pure partial model, denoted by $\mathrm{Q}^{\prime}=(\{N_{b|y}\},\{|\phi_{a|x}\rangle\})$.
Symmetrically, given a bipartite model one can consider a (pure) partial model on $\mathcal{H}_{A}$ by tracing out subsystem $B$. However, because our motivation is the compiled setting, we will focus on the partial models on $\mathcal{H}_{B}$. Furthermore, we remark that the notion of pure partial models is not vacuous. In particular, the optimal bipartite model for the CHSH inequality has a pure partial model on $\mathcal{H}_{B}$ [CHSH69]. With the notion of a partial quantum model, we define the notion of a partial (or one-sided) self-test for a bipartite model.
Definition 4.3 (Partial self-test).
The inequality $I\leq\eta^{\mathrm{Q}}$ is a partial self-test for a bipartite model $\mathrm{Q}^{*}=\left(\{P_{a|x}\},\{Q_{b|y}\},|\phi\rangle\right)$ with a pure partial model $\left(\{Q_{b|y}\},\{|\phi_{a|x}\rangle\}\right)$ if there exists a non-negative function $f(\epsilon)$ such that $f(\epsilon)\to 0$ as $\epsilon\to 0$, such that for any partial quantum model $\mathrm{Q}=\left(\{N_{b|y}\},\{\rho_{a|x}\}\right)$ achieving $I\geq\eta^{\mathrm{Q}}-\epsilon$ for $\epsilon\geq 0$, there exist a Hilbert space $\mathcal{H}_{\mathrm{aux}}$, a collection of auxiliary states $\{\sigma_{a|x}\}$ and an isometry $V:\mathcal{H}_{B}\to\mathbb{C}^{d}\otimes\mathcal{H}_{\mathrm{aux}}$ such that the following is satisfied for all $x,y,a,b$:
$\big\|VN_{b|y}\rho_{a|x}N_{b|y}V^{\dagger}-Q_{b|y}|\phi_{a|x}\rangle\!\langle\phi_{a|x}|Q_{b|y}\otimes\sigma_{a|x}\big\|_{2}$ $\leq f(\epsilon)$
$\text{and}\quad\big\|V\rho_{a|x}V^{\dagger}-|\phi_{a|x}\rangle\!\langle\phi_{a|x}|\otimes\sigma_{a|x}\big\|_{2}$ $\leq f(\epsilon),\hskip 5.69046pt$
Give the symmetry of $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ in the bipartite case, one can define a notion of partial self-test for either subsystem. Given a bipartite self-test, one can check that tracing out either subsystem results in a partial self-test. We leave it as an open question as to whether a partial self-test (say over $\mathcal{H}_{A}$ and over $\mathcal{H}_{B}$) implies that the correlation is a bipartite self-test.
4.1 Compiled self-tests from partial models
There are two main difficulties with self-testing in the compiled setting. Firstly, the correctness with respect to auxiliary systems property of the compiler (see Property (1) in Definition A.2) only guarantees that a QPT prover can prepare states (possibly mixed) $\rho_{a|x}$ over $\mathcal{H}_{B}$ that are negligible in trace distance from the post measurement states $P_{a|x}|\Psi\rangle\!\langle\Psi|P_{a|x}/p(a|x)$ of the ideal bipartite model $\mathrm{Q}$. This puts a fundamental constraint on our ability to exactly describe the set of ideal models in the compiled setting. Secondly, unlike in the nonlocal setting, it is not clear how to extract information about the measurements and states in the first round due to the homomorphic evaluation of the measurements and preparation of the states. To address these challenges we introduce the compiled counter-part of a partial quantum model.
Recall that a compiled (quantum) model $\widetilde{\mathrm{Q}}$ consists of a family of post-measurement states for “Alice” $|\widetilde{\phi}_{\alpha|\chi}\rangle$, which correspond to the state of the device following the encrypted question $\chi$, and encrypted answer $\alpha$, and a POVM $\{N_{b|y}\}$ employed by “Bob”. One could also consider a more general compiled quantum model, which includes a description of the initial state and Alice’s operators. The point of taking the coarser model is that it allows us to introduce the notion of the compiled-counterpart of a bipartite model $\mathrm{Q}$, which relates the post-measurement information in the bipartite setting with another bipartite model that resembles a compiled model.
Definition 4.4 (Compiled-counterpart model).
Given a pure partial model $\mathrm{Q}^{\prime}$, the compiled-counterpart model of $\mathrm{Q}^{\prime}$ is the pure partial model $\widetilde{\mathrm{Q}}^{(\lambda)}=(\{|\widetilde{\phi}_{\alpha|\chi}^{(\lambda)}\rangle\},\{Q_{b|y}^{(\lambda)}\})$ satisfying the following conditions for all $\lambda\in\mathbb{N}$:
$|\widetilde{\phi}_{\alpha|\chi}^{(\lambda)}\rangle$ $=|\phi_{a|x}\rangle,\ \text{ for all }\ \mathsf{sk}:\mathsf{Gen}(1^{\lambda})=\mathsf{sk},\ \chi:\mathsf{Enc}(x,\mathsf{sk})=\chi,\ \alpha:\mathsf{Dec}(\alpha,\mathsf{sk})=a.$
$N_{b|y}^{(\lambda)}$ $=Q_{b|y},\ \text{ for all }b,y.$
We remark that the compiled counterpart need not be an actual compiled model. For example, it is not required to satisfy the QPT conditions needed of a compiled model. Instead it is a model that resembles an idealized version of an honest implementation of a partial model under homomorphic encryption. We proceed with a definition of self-testing in the compiled setting that resembles partial self-testing in the bipartite setting in the context of these compiled-counterparts.
Definition 4.5 (Compiled self-test).
Let $I$ denote a Bell expression with an optimal pure partial model $\mathrm{Q}^{*}$. The inequality $I\leq\eta^{\mathrm{Q}}$ is a compiled self-test for the corresponding compiled-counterpart $\widetilde{\mathrm{Q}}^{*}=(\{|\widetilde{\phi}_{\alpha|\chi}\rangle\},\{Q_{b|y}\})$, if there exists a non-negative function $f(\epsilon)$ such that $f(\epsilon)\to 0$ as $\epsilon\to 0$, such that for every pure and projective compiled model $\widetilde{\mathrm{Q}}=\left(\{|\Psi_{\alpha|\chi}\rangle\},\{N_{b|y}\}\right)$ that achieves $I\geq\eta^{\mathrm{Q}}-\epsilon$ for some $\epsilon\geq 0$, there exists a negligible function $\mathrm{negl}(\lambda)$, an isometry $V:\tilde{\mathcal{H}}\to\mathbb{C}^{d}\otimes\mathcal{H}_{\mathrm{aux}}$, and auxiliary states $|\mathsf{aux}_{\alpha|\chi}\rangle\in\mathcal{H}_{\mathrm{aux}}$, which satisfy the following for all $x,b,y$:
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big\|V|\Psi_{\alpha|\chi}\rangle-|\widetilde{\phi}_{\alpha|\chi}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big\|^{2}\leq\mathrm{negl}(\lambda)+f(\epsilon),\hskip 5.69046pt\text{and}$ (4.2a)
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big\|VN_{b|y}|\Psi_{\alpha|\chi}\rangle-Q_{b|y}|\widetilde{\phi}_{\alpha|\chi}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big\|^{2}\leq\mathrm{negl}(\lambda)+f(\epsilon).$ (4.2b)
Equation (4.2a) is a statement about the provers state after the first round. It asserts that, given a question $x$ and answer $a$, the post-measurement state is negligibly close to that of an ideal prover implementing the honest bipartite model. To see this concretely, suppose the right hand side was exactly equal to zero. Then we have the equality $V|\Psi_{\alpha|\chi}\rangle=|\widetilde{\phi}_{\alpha|\chi}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle$ for all $\chi$ such that $\mathsf{Enc}(x)=\chi$ and all $\alpha$. Substituting $|\widetilde{\phi}_{\alpha|\chi}\rangle$ for the states $|\phi_{a|x}\rangle$ from Definition 4.4, we obtain
$V|\Psi_{\alpha|\chi}\rangle=|\phi_{a|x}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle$ (4.3)
whenever $\mathsf{Enc}(x)=\chi$ and $\mathsf{Dec}(\alpha)=a$. That is, the post-measurement states are equal to the target states up an isometry. Therefore, we interpret (4.2a) as an approximate version of Eq. 4.3, which accounts for a finite size security parameter $\lambda$ and small errors in the Bell violation $\epsilon$. Equation (4.2b) is the analogous statement including the measurements in the second round. We remark that if $V$ could depend on the question $x$ and answer $a$, (4.2a) would trivially hold regardless of the compiled Bell violation, since the states $|\phi_{a|x}\rangle$ could be prepared directly. It is therefore essential to enforce the same isometry is applied for all $a$ and $x$. Furthermore, (4.2b) captures several existing self-testing results in the compiled setting. For example those presented in [NZ23, Lemma 34][CMM+24, Theorem 3.6] and [BVB+24, Eqs. 98 and 103]. Our proposed definition then goes further by also certifying the states after the first round but before Bob’s measurements, as captured by (4.2a).
It is natural to ask if Definition 4.5 is the strongest form of self-testing possible in this scenario, or if one can also certify the initial state $|\Psi\rangle$ before Alice’s measurements. An initial guess would be to show there exists an isometry $V$ satisfying
$V|\Psi\rangle\approx_{\mathrm{negl}(\lambda)}|\phi\rangle\otimes|\mathsf{aux}\rangle,$ (4.4)
where $|\phi\rangle$ is the ideal bipartite entangled state. However, on its own this statement is not very useful: such an isometry always exists, namely, one which ignores $|\Psi\rangle$ and prepares $|\phi\rangle$ directly. A possible way around is to demand the same $V$ also satisfies (4.2). At a glance, this suggests certifying the initial state alone is not meaningful in the single prover setting; one always needs to also consider the measurements. This contrasts the two prover setting, where self-testing statements made only about the state are known [ŠB20] and non-trivial due to the space-like separation of the provers. Another question worth asking is if the assumption of having a pure projective models $\widetilde{\mathrm{Q}}$ can be relaxed in the definition Definition 4.5.
@@ 4.2 Compiled self-test for tilted-CHSH inequalities
Our final result is that the extended tilted-CHSH Bell inequalities are compiled self-tests according to Definition 4.5. In particular, we have the following result.
@@
Let $\theta\in(0,\pi/4]$, $\phi\in\big(\max\{-2\theta,-\pi+2\theta\},\min\{2\theta,\pi-2\theta\}\big)\setminus\{0\}$, and let $I_{\theta,\phi}$ be the generalized tilted-CHSH functional with quantum bound $\eta^{\mathrm{Q}}_{\theta,\phi}$ according to Eq. 3.15. Then the inequality $I_{\theta,\phi}\leq\eta^{\mathrm{Q}}_{\theta,\phi}$ is a compiled self-test for the compiled-counter part of Eq. 3.18 according to Definition 4.5.
The proof is reminiscent of the approach in [BP15], and includes similar calculations to those used in [NZ23, BVB+24] which establish rigidity statements in the compiled setting. The remainder of this section is devoted to proving the above theorem. Recall the SOS decomposition for $\bar{S}_{\theta,\phi}$, given by the polynomials $N_{0},N_{1}$ defined in Eq. 3.16. Then for compiled models $I_{\theta,\phi}=\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\geq\eta^{\mathrm{Q}}_{\theta,\phi}-\epsilon$, for some $\epsilon\geq 0$, implies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]=\eta^{\mathrm{Q}}_{\theta,\phi}-\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\leq\epsilon$. Using the calculation in Eq. A.7 with $P=N_{0}$ and $P=N_{1}$ respectively, we find
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]$ $\approx_{\mathrm{negl}(\lambda)}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\Big|(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\Big|^{2}|\Psi_{\alpha|\chi}\rangle=:\hat{E}_{0},$ (4.5)
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}]$ $\approx_{\mathrm{negl}(\lambda)}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=1)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\Big|\mathbbm{1}-(-1)^{\mathsf{Dec}(\alpha)}\sin(2\theta)X_{B}-\cos(2\theta)Z_{B}\Big|^{2}|\Psi_{\alpha|\chi}\rangle=:\hat{E}_{1},$
where we defined $Z_{B}:=(B_{0}+B_{1})/(2\cos(\phi))$ and $X_{B}:=(B_{0}-B_{1})/(2\sin(\phi))$, and note that $\hat{E}_{i}\geq 0$ for $i=0,1$. This implies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{i}^{\dagger}N_{i}]\geq\hat{E}_{i}-\mathrm{negl}(\lambda)$, from which we observe that
$\hat{E}_{0}+\tau^{2}_{\theta,\phi}\hat{E}_{1}\leq\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{0}^{\dagger}N_{0}]+\tau^{2}_{\theta,\phi}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[N_{1}^{\dagger}N_{1}]+\mathrm{negl}(\lambda)^{\prime}=\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{S}_{\theta,\phi}]+\mathrm{negl}(\lambda)^{\prime}\leq\epsilon+\mathrm{negl}(\lambda)^{\prime},$ (4.6)
where $\mathrm{negl}(\lambda)^{\prime}=(1+\tau^{2}_{\theta,\phi})\mathrm{negl}(\lambda)$. Since $\hat{E}_{i}\geq 0$, we arrive at
$\hat{E}_{0}\leq\epsilon+\mathrm{negl}(\lambda)^{\prime}\ \ \mathrm{and}\ \ \tau^{2}_{\theta,\phi}\hat{E}_{1}\leq\epsilon+\mathrm{negl}(\lambda)^{\prime}.$ (4.7)
@@
Extracting the states
We begin by studying the relations imposed by $\hat{E}_{0}\leq\epsilon+\mathrm{negl}(\lambda)^{\prime}$. Define $\delta_{0}:=\epsilon+\mathrm{negl}(\lambda)^{\prime}$. We begin with several claims:
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big\|\big(\mathbbm{1}-(Z_{B})^{2}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}\leq\delta_{0}\big(1+1/\cos(\phi)\big)^{2}=:\delta_{1}$.
@@
$\big\|\big(\mathbbm{1}-(Z_{B})^{2}\big)|\Psi_{\alpha|\chi}\rangle\big\|$ $\leq\big\|Z_{B}\big((-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|$
$\ \ \ \ +\big\|\big((-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|$
$\leq(1+\|Z_{B}\|_{\mathrm{op}})\big\|\big((-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}-Z_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|.$
In the first inequality, we used the triangle inequality, and for the second inequality, we used the definition of the operator norm $\|A|\psi\rangle\|\leq\|A\|_{\mathrm{op}}\||\psi\rangle\|,\ \ \forall\;|\psi\rangle\in\mathcal{H}$. By the definition of $Z_{B}$, $\|Z_{B}\|_{\mathrm{op}}\leq(1/(2\cos(\phi)))(\|B_{0}\|_{\mathrm{op}}+\|B_{1}\|_{\mathrm{op}})=1/\cos(\phi)$. Squaring both sides, taking the expectation over $\chi:\mathsf{Enc}(x=0)=\chi$ and summing over $\alpha$, we can apply 1 to get the desired bound. ∎
@@
where we used the same steps in 5 for the first equality, and the operator inequality $\mathbbm{1}+|X_{B}|\geq\mathbbm{1}$ for the inequality. Notice that $|X_{B}|^{2}=(X_{B})^{2}$, which satisfies $\cos^{2}(\phi)(Z_{B})^{2}+\sin^{2}(\phi)(X_{B})^{2}=\mathbbm{1}$ by definition. Inserting this into the above expression we find that
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big\|\big(P_{B}^{b}-\delta_{\mathsf{Dec}(\alpha),b}\mathbbm{1}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}\leq\delta_{0}$.
@@
$\|\big(P_{B}^{b}-\delta_{\mathsf{Dec}(\alpha),b}\mathbbm{1}\big)|\Psi_{\alpha|\chi}\rangle\big\|$ $=\frac{1}{2}\|\big(\tilde{Z}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)|\Psi_{\alpha|\chi}\rangle\big\|,$
@@
$\|\big(\tilde{Z}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)|\Psi_{\alpha|\chi}\rangle\big\|\leq\|\big(Z_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)|\Psi_{\alpha|\chi}\rangle\big\|+\|\big(\tilde{Z}_{B}-Z_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|.$
Next, we square both sides of the above, and use the fact that for two real numbers $a,b$, $(a+b)^{2}\leq 2(a^{2}+b^{2})$, to obtain
$\|\big(\tilde{Z}_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}\leq 2\|\big(Z_{B}-(-1)^{\mathsf{Dec}(\alpha)}\mathbbm{1}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}+2\|\big(\tilde{Z}_{B}-Z_{B}\big)|\Psi_{\alpha|\chi}\rangle\big\|^{2}.$ (4.8)
By taking the expectation over $\chi\,:\,\mathsf{Enc}(x=0)=\chi$ and summing over $\alpha$, the first term can be bounded by applying 1, while the second can be bounded by applying 5. ∎
@@
This isometry is based on the “SWAP gate isometry” [MYS12, YN13].
Using 7, we establish the following lemma:
@@
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x=0)=\chi}\sum_{\alpha}\big\|V|\Psi_{\alpha|\chi}\rangle-|\mathsf{Dec}(\alpha)\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big\|^{2}\leq 2\delta_{0}.$ (4.10)